one number is eight more than a second number. six times the first is 8 less than 5 times the second.
Answer by mananth(16946)

let one number be x
the other number is 8 more so it will be x+8
..
6x=5(x+8)-8
6x=5x+40-8
6x=5x+32
-5x
x=32
the other number is x+8 = 32+8 =40
..
m.ananth@hotmail.ca